OpenStudy (anonymous):

solve the equation (5x+1) (x-2)=3

OpenStudy (anonymous):

(5x+1)(x-2)=3 5x^2-10x+x-2=3 5x^2-9x-2=3 5x^2-9x-5=0 did you need to find what x equals to or just up to this equation?

OpenStudy (anonymous):

Yes what milan_tiffany got

OpenStudy (anonymous):

i was writiting it out when i seen you already did :)

OpenStudy (anonymous):

what x equal to

OpenStudy (anonymous):

are you asking me

OpenStudy (anonymous):

you have to use another equation to solve for x because factoring won't work. its the equation that is -b +or- square root of b^2-4ac divided by 2(a). I'll draw it out so you can see

OpenStudy (anonymous):